# Gravity as a topological gauge theory

### v1

@@ 2.3 Associated action
where $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{6k}{\ell^{2}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{6k}{\ell^{2}}-\dfrac{2\pi^{2}\ell^{2}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ is the effective cosmological constant, and $\Lambda_{\text{vac}}$ is the vacuum contribution. Let $\Lambda_{exp}$ be the measured value of the effective cosmological constant $\Lambda$ and several situations occur according to the theoretical predictions for $\rho_{\text{vac}}$.
  • If $\Lambda_{\text{vac}}>\Lambda_{exp}$, then, agreement with experiment could be attained by fixing the appropriate value of $\Lambda_{0}=-\dfrac{6k}{\ell^{2}}$ with $k>0$, in particular, for $k=1$ $\Leftrightarrow G=SO(3,2)$ (AdS spacetime), thus potentially allowing the use of AdS/CFT correspondence [1].
  • If $\Lambda_{\text{vac}}<\Lambda_{exp}$, agreement with $\Lambda_{exp}$ would require a mutation parameter $k<0$. A particular case is $k=-1$ $\Leftrightarrow G=SO(4,1)$ (dS spacetime). Especially, $\rho_{\text{vac}}=0$ whenever we impose normal ordering as mentionned in [5].
  • The last case is $\Lambda_{\text{vac}}=\Lambda_{exp}$ this implies $k=0\Leftrightarrow G=ISO(3,1)$, which corresponds to Minkoswki spacetime. However, for $k=0$ we do not retrieve Einstein’s equations since the action consists solely of the total derivatives corresponding to both the Euler and Pontryagin densities of the curvature $R$.
@@
The equations of motion are the same as those of Einstein-Cartan gravity modulo the boundary terms if $r=0$. However, even if they do not affect the classical equations of motion, the Euler, Pontrjagin and Nieh-Yan densities (which are not present in usual Einstein-Cartan gravity) could have non trivial effects in a path integral formulation of quantum gravity when summing over all possible configurations as stated in [14, p2].
@@
For $r\neq 0$ we can identify $\alpha_{5}=-\alpha_{2}=-\dfrac{rk}{4\pi^{2}\ell^{2}}=-\dfrac{2\alpha_{1}}{3\gamma}$, with $\gamma$ the Barbero-Immirzi parameter. By replacing $e$, $r$ and $\ell$ in the action we obtain:
@@
$S_{G}[\varpi]$ $=\dfrac{3\pi}{2\Lambda_{0}G}(\mathdutchcal{E}+\dfrac{2}{3\gamma}\mathdutchcal{P})$ (15)
@@
$=\int_{\mathscr{M}}\biggl{(}\overbrace{\alpha_{1}(R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{2}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Holst}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
@@
$\qquad\qquad+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{\underbrace{\dfrac{2\alpha_{1}}{\Lambda_{0}\gamma}}_{\vartheta}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}$
@@
Thus, by building an action from a linear combination of the Euler and Pontrjagin numbers associated to a manifold with Lorentzian Cartan geometry we retrieve the Palatini action with a bare cosmological constant $\Lambda_{0}$ as well as the Euler, Pontrjagin and Nieh-Yan densities of the curvature $R$ and torsion $T$. According to [6], these three last terms should be present in any quantum theory of gravity. The main difference when comparing with the litterature is that in our case the coupling constants of the Euler and Pontrjagin densities are determined by the cosmological constant and by respectively the coupling constant of the Palatini action and the Barbero-Immirzi parameter. These dependencies on $G,\gamma$, and $\Lambda_{0}$ are interesting since a quantum version of the theory may yield effects sensitive to the bare cosmological constant $\Lambda_{0}$ in particular. We believe that probing the bare cosmological constant this way could be highly beneficial in constraining the vacuum contribution to the cosmological constant $\Lambda_{\text{vac}}$ since the two cases $\Lambda_{\text{vac}}=0$ and $\Lambda_{\text{vac}}=8\pi G\rho_{vac}$ bring about huge differences in the numerical value of $\Lambda_{0}$. One can also see that the coupling constant of the Euler density $\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}$ obtained via the equation of motion (9) for the tetrad $\beta$ matches perfectly with what has been described in [2, section 3]. And in case $r=0$, it makes the action invariant under the duality symmetry consisting of swapping $R^{ab}\leftrightarrow\tfrac{\Lambda_{0}}{3}\beta^{a}\wedge\beta^{b}$. The case of a dynamical $\ell$ will be studied in a forthcoming paper. Also, in our case, properties of Characteristic classes imply that the choice of connection has no impact on the gauge action up to exact terms. We can consider $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{6k}{\Lambda_{0}}$ as fixed. Therefore, from the way we built the action, $r$ is the only remaining free parameter and is linked to the Barbero-Immirzi parameter $\gamma$ via the relation $r=\dfrac{\pi}{\Lambda_{0}G\gamma}$.
@@
The Pontrjagin density and other terms described here correspond to the Chern-Simons modified gravity [3] terms with a constant scalar field $\vartheta=\dfrac{2\alpha_{1}}{\Lambda_{0}\gamma}$ and potential term for $\vartheta$, $S_{\vartheta}=\int_{\mathscr{M}}V(\vartheta)$, with :
$V(\vartheta)$ $=\dfrac{2\alpha_{1}}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\varepsilon_{abcd}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}}^{Cosmological~{}constant}$ (16)
$\qquad+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{i}\wedge\beta_{j}\bigr{)}}^{Nieh-Yan}.$
The Chern-Simons term also occurs in Loop Quantum Gravity (LGQ) and String theory [3].
@@
$\left\{\begin{array}[]{ll}k=-1&\Leftrightarrow G=SO(4,1)\text{ for }\Lambda_{0}>0\\[5.69054pt] \left\{\begin{array}[]{ll}k=\pm 1\Leftrightarrow G=SO(4,1),SO(3,2)\\ \ell\rightarrow\infty\\ r=0\end{array}\right\}\Leftrightarrow\left\{\begin{array}[]{ll}\Lambda_{0}\rightarrow 0\\ \gamma\rightarrow\infty\end{array}\right\}&\simeq\text{Einstein-Cartan}\\ k=1&\Leftrightarrow G=SO(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$ (25)
Taking $r\neq 0$ in the action (7) (with Pontrjagin number $\mathdutchcal{P}$ relative to the Cartan geometry $G/H$) secures the presence of the Barbero-Immirzi part of the action, together with the Pontrjagin density of $R$ and the Nieh-Yan term.
@@ 3 Characteristic polynomials and action for the Möbius group $G=SO(4,2)/\{\pm I\}$
$H\ni h=\overbrace{\begin{pmatrix}z&0&0\\ 0&\Lambda&0\\ 0&0&z^{-1}\end{pmatrix}}^{\in G_{0}=CO(4,2)}\overbrace{\begin{pmatrix}1&\bar{u}&\tfrac{1}{2}\bar{u}u\\ 0&\mathds{1}&u\\ 0&0&1\end{pmatrix}}^{\in G_{1}=\mathbb{R}^{*3,1}}\text{ where }\Lambda\in SO(3,1),\Lambda^{\text{T}}\eta\Lambda=\eta\;;\;z\Lambda\in CO(4,2)$
$u\in\mathbb{R}^{3,1}\;;\;\bar{u}=u^{\text{T}}\eta\in\mathbb{R}^{*3,1}\text{ and }z\in\mathbb{R}.$
Its Lie algebra is $\mathfrak{g}=\mathfrak{so}(4,2)=\mathfrak{g}_{-1}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$ with $\mathfrak{g}_{-1}\simeq\mathbb{R}^{3,1}$, $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)=\mathfrak{so}(3,1)\oplus\mathbb{R}$ and $\mathfrak{g}_{1}\simeq\mathbb{R}^{*3,1}$, turns out to be a graded Lie algebra [9], namely,
@@
The generators of the above graded algebra are the same as the ones described previously in (2.1) except for $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)$ which is generated by the $\{J_{ab}\}$’s and $\mathds{1}_{4}$. An element of $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)$ is $\gamma_{\mathfrak{co}}=\frac{1}{2}\gamma_{\mathfrak{so}}^{ab}J_{ab}-z\mathds{1}_{4}$ with $\gamma_{\mathfrak{so}}\in\mathfrak{so}(3,,1)$.
3.1 The Möbius group as a symmetric space
The graded algebra $\mathfrak{g}=\overbrace{\mathfrak{g}_{1}\oplus\mathfrak{g}_{0}}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{g}_{-1}}^{\mathfrak{p}}\simeq\mathbb{R}^{*3,1}\oplus\overbrace{\mathfrak{co}(4,2)}^{\mathfrak{so}(3,1)\oplus\mathbb{R}}\oplus\mathbb{R}^{3,1}$ which is reductive, is also a symmetric space $\mathfrak{g}=\mathfrak{co}(4,2)\oplus\mathfrak{m}$ with $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}\simeq\mathbb{R}^{*3,1}\oplus\mathbb{R}^{3,1}$ with commutation relations, see (LABEL:eq:gradedLa) and [10]:
@@ 3.1 The Möbius group as a symmetric space
Let $\varphi=\varphi_{\mathfrak{g}_{1}}\oplus\varphi_{\mathfrak{g}_{0}}\oplus\varphi_{\mathfrak{g}_{-1}}=\bar{a}\oplus(c-k\mathds{1}_{4})\oplus b\in\mathfrak{g}$ with $a,b\in\mathbb{R}^{3,1}$, $c\in\mathfrak{so}(3,1)$, $k\in\mathbb{R}$ and $\bar{a}=\eta a^{T}$; $\bar{b}=\eta b^{T}\in\mathbb{R}^{*3,1}$. A matrix representation for $\mathfrak{g}$ is given by
$\varphi=\begin{pmatrix}k&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-k\end{pmatrix}=\begin{pmatrix}0&\bar{a}&0\\ 0&0&a\\ 0&0&0\end{pmatrix}+\begin{pmatrix}k&0&0\\ 0&c&0\\ 0&0&-k\end{pmatrix}+\begin{pmatrix}0&0&0\\ b&0&0\\ 0&\bar{b}&0\end{pmatrix}$ (14)
and the Lie bracket on $\mathfrak{g}$ is given by the commutator:
$[\varphi,\varphi^{\prime}]$ $=\begin{bmatrix}\begin{pmatrix}k&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-k\end{pmatrix},\begin{pmatrix}k^{\prime}&\bar{a}^{\prime}&0\\ b^{\prime}&c^{\prime}&a^{\prime}\\ 0&\bar{b}^{\prime}&-k^{\prime}\end{pmatrix}\end{bmatrix}$ (15)
$=\begin{pmatrix}\bar{a}b^{\prime}-\bar{a}^{\prime}b&k\bar{a}^{\prime}-k^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c&\bar{a}a^{\prime}-\bar{a}^{\prime}a\\ k^{\prime}b-kb^{\prime}+cb^{\prime}-c^{\prime}b&[c,c^{\prime}]+b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b}&ca^{\prime}-c^{\prime}a-k^{\prime}a+ka^{\prime}\\ \bar{b}b^{\prime}-\bar{b}^{\prime}b&\bar{b}c^{\prime}-\bar{b}^{\prime}c-k\bar{b}^{\prime}+k^{\prime}\bar{b}&\bar{b}a^{\prime}-\bar{b}^{\prime}a\end{pmatrix}$
$=\begin{pmatrix}\bar{a}b^{\prime}-\bar{a}^{\prime}b&k\bar{a}^{\prime}-k^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c&0\\ k^{\prime}b-kb^{\prime}+cb^{\prime}-c^{\prime}b&[c,c^{\prime}]+b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b}&ca^{\prime}-c^{\prime}a-k^{\prime}a+ka^{\prime}\\ 0&\bar{b}c^{\prime}-\bar{b}^{\prime}c-k\bar{b}^{\prime}+k^{\prime}\bar{b}&\bar{b}a^{\prime}-\bar{b}^{\prime}a\end{pmatrix}$
A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can be decomposed as :
$\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}=\overbrace{\varpi_{1}\oplus\varpi_{0}}^{\varpi_{\mathfrak{h}}}\oplus\varpi_{-1}=$ $\varpi_{1}^{a}\begin{pmatrix}0&\bar{e}_{a}&0\\ 0&0&e_{a}\\ 0&0&0\end{pmatrix}+\begin{pmatrix}\lambda&0&0\\ 0&A&0\\ 0&0&-\lambda\end{pmatrix}+\varpi_{-1}^{a}\begin{pmatrix}0&0&0\\ e_{a}&0&0\\ 0&\bar{e}_{a}&0\end{pmatrix}$ (25)

### v3

@@ 2.3 Associated action
where $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{6k}{\ell^{2}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{6k}{\ell^{2}}-\dfrac{2\pi^{2}\ell^{2}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ is the effective cosmological constant, and $\Lambda_{\text{vac}}$ is the vacuum contribution.
Let $\Lambda_{exp}$ be the measured value of the effective cosmological constant $\Lambda$. Several situations may occur depending on the theoretical predictions for $\rho_{\text{vac}}$.
  • If $\Lambda_{\text{vac}}>\Lambda_{exp}$, then, agreement with experiment could be attained by fixing the appropriate value of $\Lambda_{0}=-\dfrac{6k}{\ell^{2}}$ with $k>0$, in particular, for $k=1$ $\Leftrightarrow G=SO(3,2)$ (AdS spacetime), thus potentially allowing the use of AdS/CFT correspondence [aharony__2008].
  • If $\Lambda_{\text{vac}}<\Lambda_{exp}$, agreement with $\Lambda_{exp}$ would require a mutation parameter $k<0$. A particular case is $k=-1$ $\Leftrightarrow G=SO(4,1)$ (dS spacetime). Especially, $\rho_{\text{vac}}=0$ whenever one imposes normal ordering as mentioned in [10].
  • The last case is $\Lambda_{\text{vac}}=\Lambda_{exp}$ this either implies $k=0\Leftrightarrow G=ISO(3,1)$, which corresponds to Minkoswki spacetime or $\Lambda_{0}=-\dfrac{6k}{\ell^{2}}\to 0$. However, for $k=0$ we do not retrieve Einstein’s equations since the action consists solely of the total derivatives corresponding to both the Euler and Pontrjagin densities of the curvature $R$.
@@
The equations of motion are the same as those of Einstein-Cartan gravity modulo the boundary terms if $r=y=0$. However, even if they do not affect the classical equations of motion, the Euler and Pontrjagin densities of $R$ (which are not present in usual Einstein-Cartan gravity) could have non trivial effects in a path integral formulation of quantum gravity when summing over all possible configurations as stated in [8, p2].
@@
Using the same relation as in [8] we can identify the Barbero-Immirzi parameter $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}=\dfrac{2e}{3r+4y}$, with our parameters. By replacing $e$, $y$ and $\ell$ in the action we obtain:
@@
$S_{G}[\varpi]$ $=\dfrac{3\pi}{2\Lambda_{0}G}\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+\dfrac{3}{4}(\dfrac{\pi}{\Lambda_{0}G\gamma}-r)\mathdutchcal{P}(\bar{\Omega}_{\mathfrak{h}})$ (16)
@@
$=\int_{\mathscr{M}}\dfrac{\alpha_{1}}{2}\biggl{(}\overbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(\dfrac{2}{\gamma}-\dfrac{2\Lambda_{0}Gr}{3\pi})R^{ab}\wedge\beta_{a}\wedge\beta_{b}}^{Holst}\overbrace{-\dfrac{\Lambda_{0}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
@@
$\qquad\qquad\overbrace{-\dfrac{3}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{\underbrace{-(\dfrac{Gr}{\pi}+\dfrac{3}{2\Lambda_{0}\gamma})}_{\tfrac{2\vartheta}{\alpha_{1}}}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{4\Lambda_{0}Gr}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}.$
Thus, by building an action from a linear combination of the Euler and Pontrjagin numbers associated to a manifold with Lorentzian Cartan geometry we retrieve the Holst action with a bare cosmological constant $\Lambda_{0}$ as well as the Euler, Pontrjagin and Nieh-Yan densities of the curvature $R$ and torsion $T$. According to [11], these three last terms should be present in any quantum theory of gravity.
@@
The main difference when comparing with the litterature is that in our case the coupling constants of the Euler and Pontrjagin densities are determined by the bare cosmological constant and by respectively the coupling constant of the Palatini action and the Barbero-Immirzi parameter as well as $r$. These dependencies on $G,\gamma$, $\Lambda_{0}$ and $r$ are interesting since a quantum version of the theory may yield effects sensitive to the bare cosmological constant $\Lambda_{0}$ in particular. We believe that probing the bare cosmological constant this way could be highly beneficial in constraining the vacuum contribution to the cosmological constant $\Lambda_{\text{vac}}$ since the two cases $\Lambda_{\text{vac}}=0$ and $\Lambda_{\text{vac}}=8\pi G\rho_{vac}$ bring about huge differences in the numerical value of $\Lambda_{0}$.
One can also see that the coupling constant of the Euler density $-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}$ obtained via the equation of motion (10) for the tetrad $\beta$ matches perfectly with what has been described in [12, section 3]. And in case $r=y=0$, it makes the action invariant under the duality symmetry consisting of the exchange $R^{ab}\leftrightarrow\dfrac{\Lambda_{0}}{3}\beta^{a}\wedge\beta^{b}$. The case of a dynamical $\ell$ will be studied in a forthcoming paper.
Also, in our case, properties of Characteristic classes imply that the choice of connection has no impact on the gauge action up to integrals of exact terms.
We can consider $e=\dfrac{4\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{6k}{\Lambda_{0}}$ as fixed. Identifying the B-I parameter as $\gamma=\dfrac{2e}{3r+4y}$ then only leaves 2 degrees of freedom corresponding to the relative value of the parameters $r$ and $y$ and to $|k|$ the absolute value of $k$.
@@
The Pontrjagin density and other terms described here correspond to the Chern-Simons modified gravity [13] terms with a constant scalar field $\vartheta=(\dfrac{Gr}{\pi}-\dfrac{3}{\Lambda_{0}\gamma})\dfrac{\alpha_{1}}{2}$ and potential term for $\vartheta$, ${\displaystyle S_{\vartheta}=\int_{\mathscr{M}}V(\vartheta)}$, with :
$V(\vartheta)$ $=(\dfrac{1}{\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})\alpha_{1}R^{ab}\wedge\beta_{a}\wedge\beta_{b}\overbrace{-\dfrac{\Lambda_{0}\alpha_{1}}{12}\varepsilon_{abcd}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}}^{Cosmological~{}constant}$ (17)
$\qquad\overbrace{-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{2\Lambda_{0}G\alpha_{1}r}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{i}\wedge\beta_{j}\bigr{)}}^{Nieh-Yan}.$
The Chern-Simons term also occurs in Loop Quantum Gravity (LQG) and String theory [13].
@@
$\left\{\begin{array}[]{ll}k=-1&\Leftrightarrow G=SO(4,1)\text{ for }\Lambda_{0}>0\\[5.69054pt] \left\{\begin{array}[]{ll}k=\pm 1\Leftrightarrow G=SO(4,1),SO(3,2)\\ \ell\rightarrow\infty\\ ¨r=y=0\end{array}\right\}\Leftrightarrow\left\{\begin{array}[]{ll}\Lambda_{0}\rightarrow 0\\ \gamma\rightarrow\infty\end{array}\right\}&\simeq\text{Einstein-Cartan}\\ k=1&\Leftrightarrow G=SO(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$ (26)
Taking $r\neq 0$ in the action (8) (with Pontrjagin number $\mathdutchcal{P}$ relative to the Cartan geometry $G/H$) secures the presence of the Barbero-Immirzi part of the action, together with the Pontrjagin density of $R$ and the Nieh-Yan term. On the other hand $y\neq 0$ is only linked to the Pontryagin and Barbero-Immirzi part of the Holst action.
@@ 3 Characteristic polynomials and action for the Möbius group $G=SO(4,2)/\{\pm I\}$
$H\ni h=\overbrace{\begin{pmatrix}\zeta&0&0\\ 0&\Lambda&0\\ 0&0&\zeta^{-1}\end{pmatrix}}^{\in G_{0}=CO(4,2)}\overbrace{\begin{pmatrix}1&\bar{u}&\tfrac{1}{2}\bar{u}u\\ 0&\mathds{1}&u\\ 0&0&1\end{pmatrix}}^{\in G_{1}=\mathbb{R}^{*3,1}}\text{ where }\Lambda\in SO(3,1),\Lambda^{\text{T}}\eta\Lambda=\eta\;;\;\zeta\Lambda\in CO(4,2)$
$u\in\mathbb{R}^{3,1}\;;\;\bar{u}=u^{\text{T}}\eta\in{\mathbb{R}^{3,1}}^{*}\text{ and }\zeta\in\mathbb{R}\setminus\{0\}.$
Its Lie algebra is $\mathfrak{g}=\mathfrak{so}(4,2)=\mathfrak{g}_{-1}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$ with $\mathfrak{g}_{-1}\simeq\mathbb{R}^{3,1}$, $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)=\mathfrak{so}(3,1)\oplus\mathbb{R}$ and $\mathfrak{g}_{1}\simeq\mathbb{R}^{*3,1}$. $\mathfrak{g}$ turns out to be a graded Lie algebra [2], namely,
@@
The generators of the above graded algebra are the same as the ones described previously in (2.1) except for $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)$ which is generated by the $\{J_{ab}\}$’s and $\mathds{1}_{4}$. An element of $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)$ is $\gamma_{\mathfrak{co}}=\frac{1}{2}\gamma_{\mathfrak{so}}^{ab}J_{ab}-z\mathds{1}_{4}$ with $\gamma_{\mathfrak{so}}\in\mathfrak{so}(3,,1)$ and $z\in\mathbb{R}$.
3.1 The Möbius algebra as a symmetric space
The graded Lie algebra $\mathfrak{g}=\overbrace{\mathfrak{g}_{1}\oplus\mathfrak{g}_{0}}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{g}_{-1}}^{\mathfrak{p}}\simeq{\mathbb{R}^{3,1}}^{*}\oplus\overbrace{\mathfrak{co}(4,2)}^{\mathfrak{so}(3,1)\oplus\mathbb{R}}\oplus\mathbb{R}^{3,1}$ of the Möbius group is reductive and is also a symmetric space $\mathfrak{g}=\mathfrak{co}(4,2)\oplus\mathfrak{m}$ where $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}\simeq{\mathbb{R}^{3,1}}^{*}\oplus\mathbb{R}^{3,1}$. The commutation relations (see (LABEL:eq:gradedLa) and [6, Example 5.2, Chapter XI]) are given by:
@@ 3.1 The Möbius algebra as a symmetric space
Let $\varphi=\varphi_{\mathfrak{g}_{1}}\oplus\varphi_{\mathfrak{g}_{0}}\oplus\varphi_{\mathfrak{g}_{-1}}=\bar{a}\oplus(c-z\mathds{1}_{4})\oplus b\in\mathfrak{g}$ with $a,b\in\mathbb{R}^{3,1}$, $c\in\mathfrak{so}(3,1)$, $z\in\mathbb{R}$ and $\bar{a}=\eta a^{T}$; $\bar{b}=\eta b^{T}\in\mathbb{R}^{*3,1}$. A matrix representation for $\mathfrak{g}$ is given by
$\varphi=\begin{pmatrix}z&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-z\end{pmatrix}=\begin{pmatrix}0&\bar{a}&0\\ 0&0&a\\ 0&0&0\end{pmatrix}+\begin{pmatrix}z&0&0\\ 0&c&0\\ 0&0&-z\end{pmatrix}+\begin{pmatrix}0&0&0\\ b&0&0\\ 0&\bar{b}&0\end{pmatrix}$ (2)
and the Lie bracket on $\mathfrak{g}$ is thus given by the commutator:
$[\varphi,\varphi^{\prime}]$ $=\begin{bmatrix}\begin{pmatrix}z&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-z\end{pmatrix},\begin{pmatrix}z^{\prime}&\bar{a}^{\prime}&0\\ b^{\prime}&c^{\prime}&a^{\prime}\\ 0&\bar{b}^{\prime}&-z^{\prime}\end{pmatrix}\end{bmatrix}$ (3)
$=\begin{pmatrix}\bar{a}b^{\prime}-\bar{a}^{\prime}b&z\bar{a}^{\prime}-z^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c&\bar{a}a^{\prime}-\bar{a}^{\prime}a\\ z^{\prime}b-zb^{\prime}+cb^{\prime}-c^{\prime}b&[c,c^{\prime}]+b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b}&ca^{\prime}-c^{\prime}a-z^{\prime}a+za^{\prime}\\ \bar{b}b^{\prime}-\bar{b}^{\prime}b&\bar{b}c^{\prime}-\bar{b}^{\prime}c-z\bar{b}^{\prime}+z^{\prime}\bar{b}&\bar{b}a^{\prime}-\bar{b}^{\prime}a\end{pmatrix}$
$=\begin{pmatrix}\bar{a}b^{\prime}-\bar{a}^{\prime}b&z\bar{a}^{\prime}-z^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c&0\\ z^{\prime}b-zb^{\prime}+cb^{\prime}-c^{\prime}b&[c,c^{\prime}]+b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b}&ca^{\prime}-c^{\prime}a-z^{\prime}a+za^{\prime}\\ 0&\bar{b}c^{\prime}-\bar{b}^{\prime}c-z\bar{b}^{\prime}+z^{\prime}\bar{b}&\bar{b}a^{\prime}-\bar{b}^{\prime}a\end{pmatrix}.$
A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can accordingly be decomposed as:
$\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}=\overbrace{\varpi_{1}\oplus\varpi_{0}}^{\varpi_{\mathfrak{h}}}\oplus\varpi_{-1}=$ $\varpi_{1}^{a}\begin{pmatrix}0&\bar{e}_{a}&0\\ 0&0&e_{a}\\ 0&0&0\end{pmatrix}+\begin{pmatrix}\lambda&0&0\\ 0&A&0\\ 0&0&-\lambda\end{pmatrix}+\varpi_{-1}^{a}\begin{pmatrix}0&0&0\\ e_{a}&0&0\\ 0&\bar{e}_{a}&0\end{pmatrix}$